For the largest battery projecxt in US, $25 million is spent on a sodium-sulfur battery storage system installed in Presideo, TX. The total power rating is four MW (mega watt). The system is composed of 80 modules and 8000 lb each. What is the power cost in $/W and power density in W/kg in this case?
1
Expert's answer
2012-09-07T11:01:44-0400
Cost C = $2.5 * 10^7, total power rating P = 4 * 10^6 W, the mass of
single
module is M = 8000 lb = 3.629 * 10^3 kg.
The power cost is: C / P = (2.5 *
10^7) / (4 * 10^6) = 6.25 $/W.
The power density is: P / (N * M) = (4 * 10^6)
/ (80 * 3.629 * 10^3) =
13.778 W/kg.